OpenStudy (anonymous):

garden shaped like rectangle. perimeter is 432 ft. lenght is 8 tmes as long as width. find length and width.

OpenStudy (anonymous):

let width be =b and then length l=8*y 2(l+b)=432 2(8b+b)=432 b=36 and l=288

OpenStudy (turingtest):

gauthami, you set up the problem right but did the algebra wrong: 2(8b+b)=432 9b=216 b=24 (width) 8b=l=192 (length) check your work: 2(192+24)=432 :)